# Evidence for Primordial Alignment: Insights from Stellar Obliquity Measurements for Compact Sub-Saturn Systems

### v1

@@ Abstract
Despite decades of effort, the mechanisms by which the spin axis of a star and the orbital axes of its planets become misaligned remain elusive. Particularly, it is of great interest whether the large spin-orbit misalignments observed are driven primarily by high-eccentricity migration — expected to have occurred for short-period, isolated planets — or reflect a more universal process that operates across systems with a variety of present-day architectures. Compact multi-planet systems offer a unique opportunity to differentiate between these competing hypotheses, as their tightly-packed configurations preclude violent dynamical histories, including high-eccentricity migration, allowing them to trace the primordial disk plane. In this context, we report measurements of the sky-projected stellar obliquity () via the Rossiter-McLaughlin effect for two sub-Saturns in multiple-transiting systems: TOI-5126 b () and TOI-5398 b (). Both are spin-orbit aligned, joining a fast-growing group of just three other compact sub-Saturn systems, all of which exhibit spin-orbit alignment. Our results strongly suggest that sub-Saturn systems are primordially aligned and become misaligned largely in the post-disk phase through violent dynamical interactions inherent to eccentric migration, as appears to be the case increasingly for other exoplanet populations.
@@ 1 Introduction
The Transiting Exoplanets Survey Satellite (TESS) mission (Ricker et al., 2015), now well into its Extended Mission phase, has produced a growing sample of sub-Saturns around bright stars. Of particular interest to this work, TESS is responsible for the discovery of at least two sub-Saturns in compact multi-planet systems: TOI-5126 b (Fairnington et al., 2023), and TOI-5398 b (Mantovan et al., 2022, 2024). Given the brightness of their host stars and advances in Extreme Precision Radial Velocity (EPRV), these systems are amenable to stellar obliquity measurements.
[…]
This paper is organized as follows. In Section 2, we outline our observational methods. In Section 3, we describe our determination of stellar parameters. In Section 4, we detail our modeling of the RM signals and, subsequently, the stellar obliquities. In Section 5, we quantify the significance of our results and discuss their implications for the origins of spin-orbit misalignment and the formation histories of sub-Saturns.
@@ 2 Observations
The RM effect of TOI-5126 and TOI-5398 were observed using the WIYN/NEID spectrograph in the High Resolution (HR) mode (, Halverson et al. 2016; Schwab et al. 2016) on February 28, 2023 and April 16, 2023, respectively. NEID is a fiber-fed (Kanodia et al., 2018), actively environmentally stabilized spectrograph (Stefansson et al., 2016; Robertson et al., 2019) with a wavelength coverage of 380 nm to 930 nm. For TOI-5126 b, we obtain 20 radial velocity (RV) measurements in HR mode with 1000-second exposures from 02:20-07:48 UT, covering its full transit. These observations occurred under atmospheric conditions with a seeing range of 1.2-2.8 (median 1.75) and an airmass range of . At a wavelength of , the NEID spectrograph achieved a signal-to-noise ratio (SNR) of 56 pixel. For TOI-5398 b, nearly the full transit (excluding ingress) was covered by 14 RV measurements obtained in HR mode with 1250-second exposures from 02:55-09:04 UT. The conditions featured a seeing range of 0.7-1.5 (median 1.0) and an airmass range of . At the same wavelength, the spectrograph’s SNR was 60 pixel. An additional 14 out-of-transit NEID observations of TOI-5398 b were obtained between April 8, 2022, and December 27, 2023, with 300-second exposure times.
The NEID spectra were analyzed using version 1.3.0 of the NEID Data Reduction Pipeline (NEID-DRP)222Detailed information is available at: https://neid.ipac.caltech.edu/docs/NEID-DRP/, and radial velocities were derived via the cross-correlation function (CCF) method within NEID-DRP. We extracted the resulting barycentric-corrected radial velocities for re-weighted orders (CCFRVMOD) from the NExScI NEID Archive333https://neid.ipac.caltech.edu/. The RV data gathered from NEID for this study are illustrated in Figure 1 (in baseline-subtracted form) and available through the Data Behind the Figure program (see figure caption).
@@ 3.1 Synthetic Spectral Fit By iSpec
We adopt the co-added NEID spectra of TOI-5126 and TOI-5398 to determine their respective stellar atmospheric parameters, including stellar effective temperature ( ), surface gravity ( ), metallicity ( ), and projected rotational velocity ( ). We use the synthetic spectral fitting technique provided by the Python package iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019) to measure these parameters.
We employ the SPECTRUM radiative transfer code (Gray & Corbally, 1994), the MARCS atmosphere model (Gustafsson et al., 2008), and the sixth version of the GES atomic line list (Heiter et al., 2021), all integrated within iSpec, to create a synthetic model for the combined NEID spectra (SNR of 206 for TOI-5126; SNR of 176 for TOI-5398). In our fitting process, we treat micro-turbulent velocities as a variable, allowing us to accurately represent the small-scale turbulent motions in the stellar atmosphere. Conversely, we determine macro-turbulent velocities using an empirical relationship (Doyle et al., 2014) that leverages established correlations with various stellar attributes. We select specific spectral regions to streamline the fitting process, focusing on the wings of the H, H, and Mg I triplet lines, which are indicative of and , as well as Fe I and Fe II lines, which are crucial for constraining and . We then utilize the Levenberg-Marquardt nonlinear least-squares fitting algorithm (Markwardt, 2009) to refine our spectroscopic parameters by iteratively minimizing the value between the synthetic and observed spectra. The final spectroscopic parameters are detailed in Table 1.
@@ 3.2 SED+MIST Fit By EXOFASTv2
As mentioned in Fairnington et al. (2023) and Mantovan et al. (2024) respectively, both TOI-5126 and TOI-5398 exhibit strong stellar rotation signals, providing additional stringent constraints on their stellar ages. We adopt Equation 7 from Barnes (2007) to calculate the gyrochronological ages for these stars based on the stellar rotation periods derived from our work (see Section 4). The resulting gyrochronological ages for TOI-5126 and TOI-5398 are Gyr and Gyr, respectively. Subsequently, we adopt these ages and their 3 uncertainties as priors for the stellar ages in our SED fit.
We perform the SED fitting using the Differential Evolution Markov Chain Monte Carlo (DEMCMC) technique, integrated within EXOFASTv2 (Eastman, 2017; Eastman et al., 2019), from which we obtain uncertainty estimates. The DEMCMC procedure was considered converged when the Gelman-Rubin diagnostic (; Gelman & Rubin, 1992) fell below 1.01 and the count of independent draws surpassed 1000. Our final stellar parameters are listed in Table 1.
| Description | Priors | TOI-5126 | Priors | TOI-5398 | |
@@ 4 Obliquity Modeling
We determine the sky-projected spin-orbit angle for TOI-5126 b and TOI-5398 b by jointly fitting their two-minute cadence TESS lightcurves and radial velocities from NEID using a modified version of allesfitter package (Günther & Daylan, 2021), along with ancillary ground-based data available from their discovery papers. The modified version of allesfitter incorporates the Hirano et al. (2011) RM model implemented in tracit (Hjorth et al., 2021; Knudstrup & Albrecht, 2022). We exclude the transit data for the non-sub-Saturn planet in both systems but model their RVs (TOI-5126 c at and TOI-5398 c at ).
@@
For TOI-5126 b, we modeled eight full transits and one partial transit from TESS Sector 45, four full transits and one partial transit from Sector 46, and four full transits from Sector 48. Within allesfitter, we simultaneously fit the TESS and NEID data together with all photometric data utilized in TOI-5126 b’s discovery paper (Fairnington et al., 2023), including observations from the CHaracterising ExOplanets Satellite (CHEOPS, Benz et al. 2021) as well as ground-based telescopes at McDonald Observatory in Fort Davis, TX and at the Cerro Tololo Inter-American Observatory (CTIO) in Cerro Tololo, Chile, which are both part of the Las Cumbres Observatory Global Network (LCOGT; Brown et al. 2013). We do not make use of the out-of-transit radial velocities presented in Fairnington et al. (2023) due to their relatively large scatter.
@@
For TOI-5398 b, we model two full transits from TESS Sector 48, as well as photometry from the LCOGT telescopes at McDonald Observatory and Teide Observatory and from the KeplerCam at the Fred Lawrence Whipple Observatory (FLWO) on Mount Hopkins, Arizona (all sourced from Mantovan et al. 2024). We additionally incorporate the out-of-transit radial velocities originally presented in Mantovan et al. (2024), which were obtained from the High Accuracy Radial velocity Planet Searcher (HARPS-N, Cosentino et al. 2012) at the Italian Telescopio Nazionale Galileo. To account for strong stellar activity in the RVs, we apply a rotational Gaussian Process (GP) regression kernel formulated by Foreman-Mackey et al. (2017),
@@
where represents the period of stellar rotation, denotes the coherence timescale, signifies the interval between two successive data points, is a scaling hyperparameter that adjusts the amplitude, and serves as the balance parameter for the periodic and nonperiodic components of the GP kernel.
@@
For both TOI-5126 b and TOI-5398 b, we adopt priors but utilize the results reported in the literature (Fairnington et al. 2023 and Mantovan et al. 2024, respectively) as our initial guesses for , , , , , and (see Table 1 for parameter definitions). We fixed eccentricity () and the argument of periastron () to 0 due to the near-zero eccentricity reported by the discovery papers of TOI-5126 b and TOI-5398 b. The two transformed quadratic limb-darkening coefficients and are initialized at 0.5 for each instrument. As an initial guess for , for which we apply Gaussian priors, we adopt the values derived from our synthetic spectral fit. Finally, we initialize at for both systems. All parameters (except and ) are allowed to vary during the fitting process. Note that, to account for the short-term overnight instrumental systematics and stellar variability in the RM fit, we employ a quadratic function.
@@
We find good agreement ( difference) with all stellar, planetary, and derived parameters reported in the discovery papers of both systems (Fairnington et al. 2023 for TOI-5126 b and Mantovan et al. 2024 for TOI-5398 b), except (2.5 discrepant) for TOI-5126, which may be due to tentative transit timing variations (see Fairnington et al. 2023). Critically, our RM fits show that both sub-Saturns are consistent with spin-orbit alignment, with TOI-5126 b having a best-fit and TOI-5398 b having a best-fit . Figure 1 reveals the best-fit joint models and associated residuals for both systems, while Table 1 contains all fitted and derived parameters.
It is useful to additionally measure the 3D stellar obliquity , which yields a system’s true spin-orbit configuration. To estimate , we first derive the stellar inclination and stellar equatorial velocity for both TOI-5126 and TOI-5398 by applying the Bayesian inference method presented by Masuda & Winn (2020) and Hjorth et al. (2021) to our values for stellar radius , stellar rotation period , and . We compute the stellar rotation periods via the autocorrelation function (ACF) implemented in SpinSpotter (Holcomb et al., 2022), finding days and days for TOI-5126 and TOI-5398, respectively. The resulting stellar rotation periods are consistent with those reported by Fairnington et al. (2023) for TOI-5126 ( days) and Mantovan et al. (2024) ( days) for TOI-5398. From our Bayesian analysis, we find and for TOI-5126 and and for TOI-5398, and compute the true 3D stellar obliquities as (see Eqn. 9 of Fabrycky & Winn 2009):
where is the inclination angle of the planet. For TOI-5126 b, we find a true obliquity of . Similarly, we derive for TOI-5398 b, demonstrating that both systems are consistent with alignment or near-alignment. Note that, in our derivation of , we adopt an uncertainty floor of 10% in the stellar rotation periods due to the latitudinal differential rotation as suggested in Epstein & Pinsonneault (2014) and Aigrain et al. (2015) and a systematic uncertainty floor of 4.2% in stellar radii reported by Tayar et al. (2022)

### v2

@@ Abstract
Despite decades of effort, the mechanisms by which the spin axis of a star and the orbital axes of its planets become misaligned remain elusive. Particularly, it is of great interest whether the large spin-orbit misalignments observed are driven primarily by high-eccentricity migration — expected to have occurred for short-period, isolated planets — or reflect a more universal process that operates across systems with a variety of present-day architectures. Compact multi-planet systems offer a unique opportunity to differentiate between these competing hypotheses, as their tightly-packed configurations preclude violent dynamical histories, including high-eccentricity migration, allowing them to trace the primordial disk plane. In this context, we report measurements of the sky-projected stellar obliquity () via the Rossiter-McLaughlin effect for two sub-Saturns in multiple-transiting systems: TOI-5126 b () and TOI-5398 b (). Both are spin-orbit aligned, joining a fast-growing group of just three other compact sub-Saturn systems, all of which exhibit spin-orbit alignment. In aggregate with archival data, our results strongly suggest that sub-Saturn systems are primordially aligned and become misaligned largely in the post-disk phase, as appears to be the case increasingly for other exoplanet populations.
@@ 1 Introduction
The Transiting Exoplanets Survey Satellite (TESS) mission (Ricker et al., 2015), now well into its Extended Mission phase, has produced a growing sample of sub-Saturns around bright stars. Of particular interest to this work, TESS is responsible for the discovery of at least two sub-Saturns in compact multi-planet systems: TOI-5126 b (Fairnington et al., 2023), and TOI-5398 b (Mantovan et al., 2022, 2024a). Given the brightness of their host stars and advances in Extreme Precision Radial Velocity (EPRV), these systems are amenable to stellar obliquity measurements.
In this study, we present Rossiter-McLaughlin (RM) effect (Holt, 1893; Rossiter, 1924; McLaughlin, 1924; Queloz et al., 2000) measurements for both confirmed compact TESS sub-Saturn systems, TOI-5126 and TOI-5398, using the high-precision NEID spectrograph (Schwab et al., 2016) on the WIYN telescope at Kitt Peak, AZ. TOI-5126, a relatively bright () late-type F dwarf (), hosts a hot sub-Saturn with a 5.5-day period and an outer warm sub-Neptune with a 17.9-day period (Fairnington et al., 2023). TOI-5398, a bright () early-type G dwarf (), harbors a warm sub-Saturn with a 10.6-day period and an inner hot sub-Neptune with a 4.8-day period (Mantovan et al., 2022, 2024a). Our reported spin-orbit angles for these two systems represent a significant contribution to the limited sample of sub-Saturns in compact multi-planet systems with such measurements. This work is additionally part of the 11th published outcome of the Stellar Obliquities in Long-period Exoplanet Systems (SOLES) survey, as detailed in a series of publications (Rice et al. 2021; Wang et al. 2022; Rice et al. 2022; Rice et al. 2023a; Hixenbaugh et al. 2023; Dong et al. 2023; Wright et al. 2023; Rice et al. 2023b; Lubin et al. 2023; Hu et al. 2024).
This paper is organized as follows. In Section 2, we outline our radial velocity observations. In Section 3, we describe our determination of stellar parameters. In Section 4, we detail our modeling of the RM signals and, subsequently, the stellar obliquities. In Section 5, we quantify the significance of our results and discuss their implications for the origins of spin-orbit misalignment and the formation histories of sub-Saturns.
@@ 2 Observations
The RM effect of TOI-5126 and TOI-5398 were observed using the WIYN/NEID spectrograph in the High Resolution (HR) mode (, Halverson et al. 2016; Schwab et al. 2016) on February 28, 2023 and April 16, 2023, respectively. NEID is a fiber-fed (Kanodia et al., 2018), actively environmentally stabilized spectrograph (Stefansson et al., 2016; Robertson et al., 2019) with a wavelength coverage of 380 nm to 930 nm. For TOI-5126 b, we obtain 20 radial velocity (RV) measurements in HR mode with 1000-second exposures from 02:20–07:48 UT. In total, 14 of these RVs cover the full transit while the remaining six provide out-of-transit baseline; one occurring pre-ingress and the other five spanning nearly post-egress. These observations occurred under atmospheric conditions with a seeing range of 1.2–2.8 (median 1.75) and an airmass range of –. At a wavelength of , the NEID spectrograph achieved a signal-to-noise ratio (SNR) of 56 pixel-1. For TOI-5398 b, we obtain 14 RV measurements in HR mode with 1250-second exposures from 02:55–09:04 UT, eight of which were in-transit, sampling nearly the entire transit (excluding ingress). The conditions featured a seeing range of 0.7–1.5 (median 1.0) and an airmass range of –. At the same wavelength, the spectrograph’s SNR was 60 pixel-1. We obtain an additional 14 NEID RVs covering TOI-5398 b’s orbit interspersed between April 8, 2022 and December 27, 2023, four with 240-second exposure times and nine with 300-second exposure times, and we utilize these measurements in subsequent global fitting for this system.
The NEID spectra were analyzed using version 1.3.0 of the NEID Data Reduction Pipeline (NEID-DRP)222Detailed information is available at: https://neid.ipac.caltech.edu/docs/NEID-DRP/, and radial velocities were derived via the cross-correlation function (CCF) method within NEID-DRP. We extracted the resulting barycentric-corrected RVs (denoted as CCFRVMOD within the NEID-DRP documentation), from the NExScI NEID Archive333https://neid.ipac.caltech.edu/. The RV data gathered from NEID for this study are illustrated in Figure 1 (note that we subtract the Keplerian baseline to more clearly demonstrate the RM effect) and are available through the Data Behind the Figure program (see figure caption for link).
@@ 3.1 Synthetic Spectral Fit By iSpec
All out-of-transit NEID spectra for TOI-5126 and TOI-5398 are corrected for RV shifts, co-added, and subsequently utilized to determine the respective stellar atmospheric parameters, including stellar effective temperature ( ), surface gravity ( ), metallicity ( ), and projected rotational velocity ( ). We use the synthetic spectral fitting technique provided by the Python package iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019) to measure these parameters.
We employ the SPECTRUM radiative transfer code (Gray & Corbally, 1994), the MARCS atmosphere model (Gustafsson et al., 2008), and the sixth version of the GES atomic line list (Heiter et al., 2021), all integrated within iSpec, to create a synthetic model for all the combined out-of-transit NEID spectra (SNR of 112 for TOI-5126; SNR of 137 for TOI-5398). In our fitting process, we treat micro-turbulent velocities as a variable, allowing us to accurately represent the small-scale turbulent motions in the stellar atmosphere. Conversely, we determine macro-turbulent velocities using an empirical relationship (Doyle et al., 2014) that leverages established correlations with various stellar attributes. We select specific spectral regions to streamline the fitting process, focusing on the wings of the H, H, and Mg I triplet lines, which are indicative of and , as well as Fe I and Fe II lines, which are crucial for constraining and . We then utilize the Levenberg-Marquardt nonlinear least-squares fitting algorithm (Markwardt, 2009) to refine our spectroscopic parameters by iteratively minimizing the value between the synthetic and observed spectra. The final spectroscopic parameters are detailed in Table 1.
@@ 3.2 SED+MIST Fit By EXOFASTv2
As mentioned in Fairnington et al. (2023) and Mantovan et al. (2024a) respectively, both TOI-5126 and TOI-5398 exhibit strong stellar rotation signals, providing additional stringent constraints on their stellar ages. We adopt Equation 7 from Barnes (2007) to calculate the gyrochronological ages for these stars based on the stellar rotation periods derived from our work (see Section 4). The resulting gyrochronological ages for TOI-5126 and TOI-5398 are Gyr and Gyr, respectively. Subsequently, we adopt these ages and their 3 uncertainties as priors for the stellar ages in our SED fit.
We perform the SED fitting using the Differential Evolution Markov Chain Monte Carlo (DEMCMC) technique, integrated within EXOFASTv2 (Eastman, 2017; Eastman et al., 2019), from which we obtain uncertainty estimates. The DEMCMC procedure was considered converged when the Gelman-Rubin diagnostic (; Gelman & Rubin, 1992) fell below 1.01 and the count of independent draws surpassed 1000. Our final adopted stellar parameters are listed in Table 1; we find all derived values are consistent (within ) with results from discovery papers (see Fairnington et al. 2023 for TOI-5126, Mantovan et al. 2024a for TOI-5398).
| Description | Priorsa | TOI-5126b | Priorsa | TOI-5398 | |
@@
| Linear limb-darkening coefficient for HARPS-N | - | - | |||
| Quadratic limb-darkening coefficient for HARPS-N | - | - | |||
@@
| Linear limb-darkening coefficient for HARPS-N | - | - | - | ||
| Quadratic limb-darkening coefficient for HARPS-N | - | - | - | ||
@@ 4 Obliquity Modeling
We determine the sky-projected spin-orbit angle for TOI-5126 b and TOI-5398 b by jointly fitting transit photometry from their respective discovery papers (including two-minute cadence TESS light curves), and all in-transit and out-of-transit RVs from our NEID observations (see Section 2), using a modified version of the allesfitter package (Günther & Daylan, 2021). The modified version of allesfitter incorporates the Hirano et al. (2011) RM model implemented in tracit (Hjorth et al., 2021; Knudstrup & Albrecht, 2022). We exclude the transit data for the non-sub-Saturn planet in both systems but model their RVs (TOI-5126 c at and TOI-5398 c at ).
@@
For TOI-5126 b, we adopt the Presearch Data Conditioning Simple Aperture Photometry (PDCSAP) light curves from the TESS Science Processing Operations Center (SPOC), modeling eight full transits and one partial transit from Sector 45, four full transits and one partial transit from Sector 46, and four full transits from Sector 48. Within allesfitter, we simultaneously fit the TESS and NEID data together with all photometric data utilized in TOI-5126 b’s discovery paper (Fairnington et al., 2023), including observations from the CHaracterising ExOplanets Satellite (CHEOPS, Benz et al. 2021) as well as ground-based telescopes at McDonald Observatory in Fort Davis, TX and at the Cerro Tololo Inter-American Observatory (CTIO) in Cerro Tololo, Chile, which are both part of the Las Cumbres Observatory Global Network (LCOGT; Brown et al. 2013). Due to their large scatter relative to the expected planetary RV semi-amplitudes, we do not make use of the out-of-transit RVs presented in Fairnington et al. (2023), which were obtained from the Tillinghast Reflector Echelle Spectrograph (TRES) at the Fred Lawrence Whipple Observatory (FLWO) and the CHIRON spectrograph on the SMARTS telescope at CTIO.
@@
For TOI-5398 b, we also consider the PDCSAP TESS light curves (we did not find substantive evidence for the over-corrections in the PDCSAP data reported by Mantovan et al. 2024a), modeling two full transits from TESS Sector 48, as well as photometry from the LCOGT telescopes at McDonald Observatory and Teide Observatory and from the KeplerCam at FLWO on Mount Hopkins, Arizona (all sourced from Mantovan et al. 2024a). We additionally incorporate the out-of-transit RVs from Mantovan et al. (2024a), which were obtained from the High Accuracy Radial velocity Planet Searcher (HARPS-N, Cosentino et al. 2012) at the Italian Telescopio Nazionale Galileo (TNG). Recently, Mantovan et al. (2024b) published an independent RM effect measurement of TOI-5398 b; therefore, we model the RM effect for this target both with and without the inclusion of their published in-transit RVs (holding all other factors equal), which were also obtained from the HARPS-N spectrograph444Our work was otherwise prepared concurrently and independently to that of Mantovan et al.. To account for strong stellar activity in the RVs, we apply a rotational Gaussian Process (GP) regression kernel formulated by Foreman-Mackey et al. (2017),
@@
where represents the period of stellar rotation, denotes the coherence timescale, signifies the interval between two successive data points, is a scaling hyperparameter that adjusts the amplitude, and serves as the balance parameter for the periodic and nonperiodic components of the GP kernel. For TOI-5398, we note that our adopted GP model is most similar to “Case 1” from Mantovan et al. (2024a), though the results from their multi-dimensional “Case 3” were ultimately adopted.
@@
For both TOI-5126 b and TOI-5398 b, we adopt priors but utilize the results reported in the literature (Fairnington et al. 2023 and Mantovan et al. 2024a, respectively) as our initial guesses for , , , , , and (see Table 1 for parameter definitions). We fixed eccentricity () and the argument of periastron () to 0 due to the near-zero eccentricity reported by the discovery papers of TOI-5126 b and TOI-5398 b. We consider transformed linear and quadratic limb-darkening coefficients and , in addition to physical linear and quadratic limb-darkening coefficients and 555The relationship between the transformed (, ) and physical (, ) limb-darkening coefficients is described in Kipping (2013) by Equation (15): and Equation (16): , for both NEID and HARPS-N, initializing and at 0.5 for each instrument. As an initial guess for , for which we apply Gaussian priors, we adopt the values derived from our synthetic spectral fit. Finally, we initialize at for both systems. All parameters (except and ) are allowed to vary during the fitting process. To account for the short-term overnight instrumental systematics and stellar variability in the RM fit, we employ a quadratic baseline for TOI-5126 and a constant baseline for TOI-5398, though we note that all tested baselines (constant, linear, quadratic, and cubic) produced consistent values (within ).
@@
We find good agreement ( difference) with all stellar, planetary, and derived parameters reported in the discovery papers of both systems (Fairnington et al. 2023 for TOI-5126 b and Mantovan et al. 2024a for TOI-5398 b), except (2.5 discrepant) for TOI-5126, which may be due to tentative transit timing variations (see Fairnington et al. 2023). Critically, our RM fits show that both sub-Saturns are consistent with spin-orbit alignment. For TOI-5126 b, we find a best-fit projected obliquity of . For TOI-5398 b, we compute a best-fit using in-transit RVs from both NEID (presented in this work) and HARPS-N (presented in Mantovan et al. 2024b), while we derive a marginally aligned value of using our NEID data alone (note that both values are within of the reported by Mantovan et al. 2024b). Figure 1 reveals the best-fit joint models and associated residuals for both systems, while Table 1 contains all fitted and derived parameters.
It is useful to additionally measure the 3D stellar obliquity , which yields a system’s true spin-orbit configuration. To estimate , we first derive the stellar inclination and stellar equatorial velocity for both TOI-5126 and TOI-5398 by applying the Bayesian inference method presented by Masuda & Winn (2020) and Hjorth et al. (2021) to our values for stellar radius , stellar rotation period , and . We utilize the aforementioned TESS PDCSAP light curves to compute the stellar rotation periods for each target, employing the autocorrelation function (ACF) implemented in SpinSpotter (Holcomb et al., 2022). Correspondingly, we find days for TOI-5126 and days for TOI-5398. However, due to the effects of latitudinal differential rotation (Epstein & Pinsonneault, 2014; Aigrain et al., 2015) and the systematic uncertainty floor (of 4.2%) on the derivation of stellar radii (Tayar et al., 2022), we adopt a 10% uncertainty floor on our derived stellar rotation periods, yielding days and days for TOI-5126 and TOI-5398, respectively. These estimates are consistent with those reported by their respective discovery papers, which both applied the Lomb-Scargle periodogram to photometric data, rather than the ACF, in order to ascertain their adopted rotation periods (Fairnington et al. 2023 find days for TOI-5126, Mantovan et al. 2024a find days for TOI-5398). From our Bayesian analysis, we find and for TOI-5126 and and for TOI-5398, and compute the true 3D stellar obliquities as (see Eqn. 9 of Fabrycky & Winn 2009):
where is the inclination angle of the planet. For TOI-5126 b, we find a true obliquity of . Similarly, we derive for TOI-5398 b (in good agreement with the derived by Mantovan et al. 2024b), demonstrating that both systems are consistent with alignment.